• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eric

    Log entry time 00:09:24 on June19,2002

    Entry number 86364

    keyword=Shift summary


    New crew on shift: Guillaume Batigne, Pierre Moussiegt, Eric Voutier, Bodo Reitz
    ***************

    16:00 Sweeping Hall A in progress
    16:43 Power permit
    17:15 Beam permit
    19:15 MCC calls, had problems with the injector and will startonly now to restore beam.... if everything goes well, we may expect beam by 20:30.
    20:30 Update... after they solve the RF problem at the injector, they experience a "phase" problem and are now steering again >> in practice they are still at the injector
    level, expect to pass it now at resume beam within 1-2 hours.
    20:57 Pulse beam to the hall
    21:29 Ready to send CW beam to the hall; MCC starts with low current to check the
    line
    21:35 Beam is on with 5 uA; start an HARP scan to check the beam spot
    Harp run number 1791 ..... coda run # 1621 ..... 101 um x 57 um
    21:50 MCC wants to try to go to 100 uA before we continue to tune the beam spot
    since further adjustement may affect the beam spot too. Carbon target being in place
    we start pointing meassurement at the same time.
    21:59 Take beam for spot adjustement, we end with 69 um x 81 um and take this
    beam spot for data taking
    22:16 Start data taking
    22:24 Move to 15 cm LD2 target
    22:33 Beam is unstable / MCC check something in our line
    23:52 Start taking data at 70 uA, the beam is stable there and they will try to increase
    increase intensity over the next hour.

    Run history
    --------

    >- Q2_f40l -<

    1622 ...... Pointing / beam unstable
    1623 ...... Pointing / short run / MCC takes beam for beam spot adjustement
    1624 ...... Pointing / final good run
    1625 ...... 15 cm LD2